OpenStudy (anonymous):

√5(10-4√2) ?????????????

OpenStudy (anonymous):

\[\Large 10\sqrt 5 - 4\sqrt {10} \]

OpenStudy (anonymous):

sqrt(50-20 sqrt(2))?

OpenStudy (anonymous):

Multiply out the brackets. \[(\sqrt{5} \times 10) - (4 \times \sqrt{2} \times \sqrt{5})\]\[=10 \sqrt{5} - 4 \sqrt{10}\]